answersLogoWhite

0

If they are fair coins, the probability is 0.25

User Avatar

Wiki User

∙ 12y ago

What else can I help you with?

Related Questions

What is the probability of flipping 3 coins and getting all heads or tails?

The probabilty of you flipping 3 coins and getting all heads or tails is 0.125 or 1/8.


What is the probability of flipping 5 coins and getting all tails?

The probability of getting all tails is 1/25 = 1/32


What is the probability of flipping three tails when three coins are flipped?

The probability of flipping three tails with three coins is (1 in 2)3 or 1 in 8 or 0.125.


If you Flip four coins at once what is probability of 2 head and 3 tail?

http://wiki.answers.com/Q/If_you_Flip_four_coins_at_once_what_is_probability_of_2_head_and_3_tail" The probability of flipping four coins and getting 2 heads and 3 tails is ZERO 2 heads and 3 tails requires flipping FIVE coins.


What is the probability of flipping 4 fair coins 1 time and getting all tails?

The probability of getting all tails when you flip 4 fair coins 1 time is:P(4 tails) = (1/2)4 = 1/16 = 0.0625 = 6.25%


What is the probability of flipping no heads when you flip three identical coins?

The probability of flipping one coin and getting tails is 1/2. In order to find the probability of multiple events occurring, you find the product of all the events. For 3 coins the probability of getting tails 3 times is 1/8 because .5 x .5 x .5 = .125 or 1/8.


What is the probability of flipping 4 coins and getting all heads or all tails?

The probability of getting all heads is 1/24 = 1/16 The probability of getting all tails is also 1/24 = 1/16 The probability of all heads or all tails is the sum of the two = 1/8


What is the probability of flipping two coins and one landing on heads and one landing on tails?

It is 1/2.


What is the probability of rolling a prime number and flipping a coin and getting tails?

1 in 2.


What is the probability of flipping a coin ten times and getting tails six times?

ye


If you flip two coins what is the probability that they will both land on tails?

When flipping two coins, each coin has two possible outcomes: heads (H) or tails (T). The total number of outcomes when flipping two coins is 2 × 2 = 4, which includes HH, HT, TH, and TT. Out of these four outcomes, only one results in both coins landing on tails (TT). Therefore, the probability of both coins landing on tails is 1 out of 4, or 25%.


What is the probability of flipping a coin and it landing on tails how many coins bill has?

The answer to the first question is 0.5. The answer to the second is not possible to work out.

Trending Questions
What does the term relative risk refer to? If you draw a card from a standard deck of cards what is the probability of drawing a heart or a face card? Eighty one percent of us households have DVD players choose 6 households at random what is the probability that at least one does not have a DVD player Please Show Work? What is average temp of a 10year old? The collection of all possible individuals objects or measurements is called? How do you increase figures by 20 percent on excel? How do you calculate deviation sensitivity? Does last couple of days mean last two days? What is a draw bar used for? What graphs are most effective at displaying data that change continuously? How much is the average weight for 5 ft 10? If the F statistic is greater than 1.0 the rejection region is in the left tail if the value is less than 1.0 the rejection region is in the right tail? How many even 3-digit numbers can be formed from the digits 1 2 5 6 9 if each digit can be used only once? The set of all for an event is called the sample space? A bag contains 5 red marbles 3 blue marbles and 2 green marbles What is the probability of pulling out two different color marbles? What is the probability of pulling the right card from a deck of cards? How many pounds is a 13 year old supposed to be? What measures responsiveness that shows how a dependent variable responds to a change in an independent variable? Identify the three things you must do to assess the validity of a poll? Is weight normally distributed?